OpenStudy (minatsukisaya95):

x=a(b-sinb),y=a(1-cos b), prove dy/dx=cot(b/2)

OpenStudy (anonymous):

dy/db = asinb dx/db = a(1-cosb) helps ?

OpenStudy (anonymous):

another hint : now you can divide dy/db by dx/db and get dy/dx !

OpenStudy (anonymous):

ill help more: ( dy/db ) / (dx/db) = asinb/ a(1-cosb) = sinb / (1-cosb) = dy/dx

OpenStudy (anonymous):

now it is just trigo.

OpenStudy (anonymous):

if you dont like the idea of dividing dy/db by dx/db you can as well use the chain rule : dy/dx = dy/db * db/dx and you can find db/dx by isolating b from the equation with x and finding db/dx.

OpenStudy (minatsukisaya95):

the answer i got is different from what to prove

OpenStudy (anonymous):

you get dy/dx = sinb / (1-cosb)

OpenStudy (anonymous):

now it can be shown that it is the same as cot(b/2)

OpenStudy (anonymous):

hint : sin(b) = sin(2 * b/2) cos(b) = cos(2*b/2) use the formulas for double angle now.

OpenStudy (minatsukisaya95):

yes i get that but how do we solve that to get cot(b/2)

OpenStudy (anonymous):

look at my last response it should help you

OpenStudy (anonymous):

sin(b) = 2sin(b/2)cos(b/2) cos(b) = cos^2(b/2) - sin^2(b/2) so sinb / (1-cosb) = 2sin(b/2)cos(b/2) / (1- cos^2(b/2) + sin^2(b/2))

OpenStudy (anonymous):

i hope you can do it from here.

OpenStudy (anonymous):

@minatsukisaya95 are you there ?

OpenStudy (minatsukisaya95):

yes thank u i got it

OpenStudy (anonymous):

yw